OpenStudy (anonymous):

How do you graph this? 4y + 17 greater than or equal to 29

OpenStudy (anonymous):

4y+17>=29 4y>=29-17 4y>=12 y>=3

OpenStudy (anonymous):

How did you get that answer tho?

OpenStudy (anonymous):

just flow the steps 4y+17>=29 4y>=29-17 4y>=12 y>=3

OpenStudy (anonymous):